OpenStudy (anonymous):

What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form.

hero (hero):

Insert each point into y = xm + b: -4 = -1m + b 4 = 1m + b Isolate b in each equation: b = m - 4 b = 4 - m set b = b m - 4 = 4 - m m + m = 4 + 4 2m = 8 m = 8/2 m = 4 b = m - 4 b = 4 - 4 b = 0 y = 4x

OpenStudy (anonymous):

thanks ! i judt didnt understand what to do haha